# Constructions are Revealed in Word Distributions

### v1

@@ 5 Generalizing to Other Constructions
Figure 4: Global affinity for CoGS. CW: Causative-with; CC: Comp-correlative; Con: Conative; LA: Let-alone; ML: Much-less; Way: Way-manner
@@
Figure 6: The affinity matrix encodes diverse types of interactions, including both constructional and non-constructional interactions.
We now investigate how far we can push this distributional approach by generalizing to two additional datasets: the Construction Grammar Schematicity corpus (CoGS, Bonial and Tayyar Madabushi 2024) and MAGPIE Haagsma et al. (2020), a large corpus of annotated potentially idiomatic expressions.
@@ 5.1 Global affinity distinguishes fixed slots in numerous constructions
CoGS contains 50 examples for each of 10 construction types (see Appendix C for examples). Six of these have at least one fixed word (partially substantive) for which we calculate global affinity. As shown in Figure 6, the fixed words—in all but the conative—have high global affinities. The at in the conative has low affinity (e.g., He kicked at the ball), which suggests that various other completions are also possible. These results show that the distribution simulated by RoBERTa captures the contextual cues associated with various partially substantive constructions, and that the global affinity method reflects that contextual affinity.
@@
5.2 Global affinity helps distinguish literal from figurative usages in potentially idiomatic expressions
@@
Next, we ask whether global affinity helps discriminate literal from figurative usages in MAGPIE, a corpus of 50,000 sentences containing potentially idiomatic expressions (PIEs) that are hand-labeled as figurative or literal usages. Idioms (e.g., kick the bucket) have long been of interest to CxG (e.g. Fillmore, 1988; Croft and Cruse, 2004; Wulff, 2008). Some idioms are entirely fixed, while others are partially-substantive and can be modified lexically or syntactically (e.g., kicked the bucket).
PIEs are of interest because, like other constructions, they may be activated by contextual semantics (e.g., Croft and Cruse 2004, p. 239 argue that interpreting the let-alone construction requires a relatively complex “semantic apparatus” that considers multiple interacting parts of the context). Here we investigate whether high global affinity (triggered by the context) may reflect figurative usage; albeit imperfectly, since other factors can constrain words (see e.g., Section 7), and since frequently occurring PIEs might have high affinities even when used literally (e.g., “nuts and bolts in the garage”; see, e.g., Rambelli et al. 2023).
To assess the signal provided by global affinity in figurative versus literal usages, we produce a receiver operating characteristic (ROC) curve, classifying by simply thresholding global affinity, and achieve an area of 0.71 under the curve (ROC plot provided in appendix, Figure 13).
6 Schematic Constructions: Covarying Slots and Abstract Categories
@@
6.1 Model distributions generalize the NPN’s noun-noun interaction
The NPN construction (e.g., “day after day”) is a schematic construction (see e.g., Sommerer and Baumann 2021 for a recent study using collostructional analysis). Since the construction is entirely schematic (no slots are constrained to be fixed words), a distributional learner can acquire it only by generalizing to abstract classes (i.e., noun and preposition).
To study whether the distribution reflects generalization of the NPN to arbitrary nouns, we randomly sample 100 singly-tokenized, singular nouns from RoBERTa’s vocabulary and prompt GPT-4 to produce four NPN sentences for each noun with each of the prepositions by, after, upon, and to,333 There are other varieties of NPN than the symmetric N+P+N and other prepositions can be used Jackendoff (2008); Sommerer and Baumann (2021); this analysis is intended only to show generalization, not to be an exhaustive study. (400 examples). After filtering to generations that match the constraints (e.g., GPT-4 struggles to produce NPNs using to; see Appendix E), we have 295 sentences. We compute global affinities for the nouns in the NPN and obtain acceptability ratings (scale 1-5) from the last author, who was blind to the affinity scores. These ratings are used only to filter out unnatural generations. In Figure 6 we plot affinity scores for nouns in sentences with acceptability $\geq$ 4 (n=162). We provide the same plot but with all sentences (any acceptability) in the appendix, Figure 14); we note that lower acceptability sentences have lower global affinity, reflecting that affinity is sensitive to linguistic acceptability.
In Figure 6, global affinity scores for upon and after suggest that the distribution captures the general form of the NPN: the model generally expects the two nouns to be the same, in spite of the fact that the sampled nouns are unlikely to have been seen as part of NPNs during training. Moreover, the lower affinity in the case of by and to accords with prior characterization of the NPN: to is only semi-productive, and after and upon are more flexible in NPN use than other prepositions Jackendoff (2008). From the counts of valid and acceptable generations (see Appendix E) we observe that GPT-4 has more difficulty producing acceptable NPNs for by and to, which is also consistent with their relative productivities.
@@ 6.2 Models generalize the comparative correlative’s category constraint
We test whether the distribution encodes the semantic category of the comparative adjective/adverb in the CC (see Weissweiler et al., 2022, for a recent study), which is of the form, e.g., The better your distribution, the more constructions it will encode. Whereas in Section 5.1, we showed that the in the CC has high global affinity, here we test whether the distribution also encodes an abstract slot constraint for the comparative adjective/adverb. Using the 54 CC examples from the CoGS dataset, we mask each comparative adjective/adverb, obtain the set of highest probability outputs at the masked position that sum to 98% probability mass, and calculate a comparative score: the percentage of this set that is a comparative adjective/adverb (see Appendix F).
Out of 99 comparative adjectives/adverbs, 95 score 100%; another 3 score $>99\%$, and one (The higher up the nicer!) scores 86%. This result shows that the distribution captures the abstract syntactic (comparative adjective/adverb) constraint of the CC nearly perfectly, with 98/99 examples having scores $\geq 99\%$.
@@ 7 The Limits of Distributional Analysis
Figure 7: Green Day (a cx) is present in top/ bottom panels but without/ with band, it has low/ high global affinity (white = 0, black = 1).
Figure 8: The affinity matrix reflects interactions (pink) between between band, Green, and Day, as expected.
@@
How far can we push these distributional approaches for identifying constructions in text? Could they allow us to identify a model’s complete constructicon bottom-up (see, e.g., Dunn, 2017, 2019, 2024; Feng et al., 2022; Lyngfelt et al., 2018; Xu et al., 2024)? Here we argue that the answer in the limit is likely “no”.
First, consider the sentence Alice went to the hardware store and bought a hammer. The affinity matrix (Figure 6) reveals various interactions of interest. Some of these are likely constructionally mediated, including tense agreement in coordinated verb phrases (went, bought), as well as subject-verb (she, bought), head-modifier (hardware, store), and determiner-noun (a, hammer) dependencies. But other interactions are less obviously constructional, such as coreference (Alice, she) and semantic relatedness (hardware, hammer). So although the affinity matrix does reflect constructional relations, it also shows that other contextual interactions can produce high local affinities. Thus affinity is insufficient 444 Insufficiency holds for global affinity as well: Figure 15 shows that hardware has high global affinity (0.92). to indicate the presence of a construction.
Second, consider the sentence My favorite band is Green Day, which includes the non-compositional collocation Green Day (a well-known band name). Figure 9 shows that Green Day has low global affinity until an appropriate contextual trigger is given, i.e., band. This shows that even substantive constructions may exhibit low global affinities when the surrounding context is insufficient to trigger them.555 See also Figure 17 in the Appendix, for an additional example with the idiom kicked the bucket.
Additionally, Figure 9 shows the affinity matrix for the Green Day example. As expected, it reflects affinity between band, Green, and Day. However, in a different context (I saw my favorite band Green Day in concert), the interactions between Green Day and band vanish (Figure 9). This appears to be due to the co-presence of an additional semantic cue (concert): since band and concert are both individually sufficient to cue the construction Green Day, neither individually exhibits strong pairwise affinity with Green Day (i.e., neither when masked substantially changes the model’s output distribution for Green or Day).
Taken together, the hardware store and Green Day examples show that affinity is neither sufficient nor inecessary for a construction to be present. Non-sufficiency arises out of the tension between contextual and constructional interactions. Non-necessity arises from a methodological challenge: eliciting global affinity requires masking, but under mask, the context may not trigger a construction. These results thus warrant caution in mapping between observables (affinities) and hypothetical constructs (like constructions). They also suggest avenues for future work, which we explore below.
@@ 9 Conclusion
Our analysis is limited to a single (bidirectional) masked language model, RoBERTa, and could be rerun on other models of different sizes. In choosing masked language models, a straightforward analysis of the sort that we have performed is also limited to sentences where the word in focus happens to be a single token. Our automatic pipeline for Section 4.3 dropped many examples, leaving few for analysis; nonetheless the approach still recovered mislabeled examples.
@@ Acknowledgements
Joshua Rozner was supported by a fellowship from the Institute for Computational and Mathematical Engineering at Stanford University. Leonie Weissweiler was supported by a postdoctoral fellowship from the German Academic Exchange Service (DAAD). We thank Shijia Zhou for providing precise results for prior work on the CEC, EAP and AAP.
@@ Appendix A Supplement: Methods
Calculating a local affinity matrix is more expensive than calculating global affinities: $n+n^{2}$ forward passes versus $n$. Note that the affinity matrix is asymmetric—positions (i,j) and (j,i) involve three different distributions and are computed as JSD(A,B) vs JSD(C,B)—and that the diagonal of the matrix, $a_{i,i}$ is 00.
We use RoBERTa large. Most experiments are run locally on an M3 Macbook Pro or on a single Nvida RTX A6000 GPU on a Stanford cluster. Total time to run all experiment code is less than an hour.
@@ B.1 Data and preprocessing
B.2 Supplement: Section 4.1: Models distinguish the CEC from the EAP and AAP in their output distributions
Figure 10: This figure is the same as Figure 3, but without correction of mislabeled examples. Note that only 5 examples differ, so the plots are very similar.
@@
B.3 Supplement: Section 4.2: Models capture causal relations in the CEC
@@ B.3.1 Dataset (multiple-that)
B.4 Supplement: Section 4.3: Affinity patterns distinguish the EAP and AAP
Figure 11: This plot shows which pairwise interactions, on average, most substantially differ between EAP and AAP examples.
Figure 12: Uncorrected multiplot version of Figure 3. UMAP projections for EAP (green) and AAP (orange). Each row corresponds to the number of dimensions that are used for the projection (2-6). Columns correspond to two different random seeds. Potentially “misclassified” sentences (those that were near the class boundary) are labeled with their original dataset IDs for discussion. No labels have been corrected in these plots.
As described in Section 4.3, we identify potentially salient differences in the EAP and AAP patterns: we compute a single average (position-wise) affinity matrix for all EAP examples and another average matrix for all AAP examples; subtract these two average matrices; and then take the position-wise absolute value. The largest values in the resulting matrix provide the interactions that seem to most distinguish the EAP and AAP. The heatmap in Figure 11 shows the most potentially informative interactions across EAP and AAP examples. We fill the diagonal with the global affinity scores as probabilities. Here we see that the most different interactions (between EAP and AAP examples) on average are that with itself, that with verb2, and verb1 with so.
After running the pre-processing pipeline for tokenization we have 68 AAP and 24 EAP examples (see Appendix B.1). For our analysis, we exclude any that have multitokenized words in one of the seven positions we consider, or for which Spacy Honnibal et al. (2020) fails to label a POS. This leaves us with 14 and 26 EAP and AAP examples, respectively, for the analysis (before any mislabeled examples are corrected; see below).
Figure 12 shows UMAP plots using different numbers of affinities: each row corresponds to the number of informative dimensions that are chosen using the heatmap. The two columns show two random seeds for the UMAP projection. From this multiplot we see that using only two or three dimensions does not separate the classes. This suggests that it is useful to allow different affinity scores to interact in a higher dimensional space during projection. On the other hand, using too many dimensions seems to include interactions that are not important to distinguishing the EAP and AAP (e.g., using six dimensions leads to less separability). Accordingly, we use five dimensions in the main text (Figure 3).
Finally, we label five sentences in the plots (Figure 12) which either look to violate the separability or which are on the boundary. We use the same three labels (misclassified, mislabeled, unclear) as were used in Section B.2 above.
@@
    Label 95 corresponds to the sentence After a series of fires…, which we identified as being unclear previously.
@@
    Label 286: Seeing this in a jar, I was so hopeful that it had it’s own well sealed jar protecting the powder from hardenin (sic)
    The sentence is originally labeled as epistemic, however we argue that it is unclear since hopeful describes affect and since there is also a causal interaction: “seeing this in a jar” leads the narrator to be “hopeful.”
@@
    Label 1: He was so afraid that rival loyalist inmates wished to kill him inside the prison.
    This sentence is originally labeled as epistemic, but afraid describes affect. This sentence is unclear since it is both epistemic and affective: The man conjectures that the “inmates wished to kill him” and this causes him to be afraid.
@@
    Label 320: I am so startled that no one seemed to have a problem with the stage being held up, the women passengers being kidnapped and sold.
    This is correctly labeled as affective and thus could be misclassified by this method depending on the random seed.
@@
    Label 270: I was so disturbed that it took me this long.
    This is correctly labeled as affective and could be misclassified by this method depending on the random seed.
Appendix C Supplement: CoGS (Section 5.1)
Examples of the the six construction types (from Bonial and Tayyar Madabushi 2024). Fixed words are italicized.
@@
    Causative-with: She loaded the truck with books.
@@
    Comparative correlative: The more the merrier. (The fixed word the occurs twice. In our analysis the two the words are considered as a single class.)
@@
    Much-less: He has not been put on trial, much less found guilty.
@@
    Way-manner: We made our way home.
Appendix D Supplement: MAGPIE (Section 5.2)
D.1 Data

### v2

@@ 5 Generalizing to Other Substantive Constructions
Figure 5: Global affinity for CoGS. CW: Causative-with; $C$: Comp-correlative; Con: Conative; LA: Let-alone; ML: Much-less; Way: Way-manner
@@
Figure 7: The local affinity matrix encodes diverse types of interactions, including both constructional and non-constructional interactions.
In this and §6, we investigate how well our distributional approaches generalize. In this section, we study generalization to partially substantive constructions using two datasets: the Construction Grammar Schematicity corpus (CoGS, Bonial and Tayyar Madabushi 2024) and MAGPIE Haagsma et al. (2020), a corpus of potentially idiomatic expressions labeled as either figurative or literal.
@@ 5.1 Global affinity distinguishes fixed slots in numerous constructions
CoGS contains roughly 50 examples for each of 10 construction types. Six of these constructions are partially substantive and have at least one fixed word for which we calculate global affinity (examples from Bonial and Tayyar Madabushi; fixed words italicized):
  • Causative-with: She loaded the truck with books.
  • Comparative correlative: The more the merrier.
    (In our analysis the two the words are considered as a single class.)
  • Conative: He kicked at the ball.
  • Much-less: He has not been put on trial, much less found guilty.
  • Way-manner: We made our way home.
As motivated in the introduction, we expect that constructions will manifest as high global affinities on fixed words. For example, just as the interactions between words in the CEC constrain so, in the let-alone cxn, we expect interactions to constrain the non-compositional let and alone.
@@
As shown in Figure 7, the fixed words—in all but the conative—have high global affinities. The at in the conative has low affinity (e.g., He kicked at the ball) because various other non-conative completions are possible: for this example the model produces out, at, over. These results show that the distribution simulated by RoBERTa captures the contextual cues associated with various partially substantive constructions, and that global affinity reflects that contextual affinity.
@@
5.2 Global affinity helps distinguish literal from figurative usages
@@
Next, we ask whether global affinity helps discriminate figurative and literal usages in potentially idiomatic expressions (PIEs). We use MAGPIE, a corpus of $\sim$50,000 sentences with PIEs that are hand-labeled as either figurative or literal. Idioms (e.g., kick the bucket) have long been of interest to CxG (e.g. Fillmore, 1988; Croft and Cruse, 2004; Wulff, 2008) and were a key motivation for the approaches in this study (see Introduction).
We hypothesize that figurative uses will have higher global affinity, by virtue of being entrenched and non-compositional: consider that one can spill the beans but that in the same context, one would not spill the water, so beans should have high affinity. Nonetheless, this signal may be confounded: other factors can constrain words (see e.g., §7), and frequent PIEs might have high affinities even when used literally (e.g., “nuts and bolts in the garage”; see, e.g., Rambelli et al. 2023).
We compute global affinity for each of 114k words that are part of a PIE in 45k sentences (10k literal, 34k figurative; details in Appendix D). Under the hypothesis, we treat affinity as a classification probability for figurative usage and produce a receiver operating characteristic curve, achieving an area of 0.71 under the curve (plot in D), which shows that global affinity provides a meaningful signal for classifying figurative vs. literal usages.333 “The AUC of a classifier is equivalent to the probability that [it] will rank a randomly chosen positive instance higher than a randomly chosen negative instance,” Fawcett (2006).
We further compare average figurative and literal scores for each idiom (graphical results in D.3). For example, nuts and bolts behaves as expected, scoring 0.90 fig and 0.84 lit, reflecting entrenchment of the literal usage. On the other hand turn someone’s head gives 0.37 vs. 0.74, but this “failure” may stem from the relative non-entrenchment of the figurative use (that Yankee sun hasn’t turned your head) versus the more common literal use (she turned her head). Moreover, qualitative analysis suggests that affinity can be useful to further characterization of PIEs: for example, the lowest scoring figurative usages are generally less familiar to the authors, and affinity helped identify some mislabeled or questionable examples. Future research combining affinity methods with other approaches might produce further quantitative insights for PIEs.
6 Generalizing to Schematic Constructions
@@
6.1 Models generalize the NPN’s covarying noun-noun slots
The NPN construction (e.g., day after day) is a schematic construction (see e.g., Sommerer and Baumann 2021 for a recent study using collostructional analysis). Since the construction is entirely schematic (no slots are constrained to be fixed words), a distributional learner can acquire it only by generalizing to abstract classes (i.e., noun and preposition).
To study whether the distribution reflects generalization of the NPN to arbitrary nouns, we randomly sample 100 singly-tokenized, singular nouns from RoBERTa’s vocabulary and prompt GPT-4 to produce NPN sentences with each of the prepositions by, after, upon, and to, [^4] There are other varieties of NPN than the symmetric N+P+N and other prepositions can be used Jackendoff (2008); Sommerer and Baumann (2021); this analysis is intended only to show generalization, not to be an exhaustive study. giving 400 sentences. We compute global affinities for the nouns in the NPN and obtain acceptability ratings (scale 1-5) from the last author, who is blind to the affinity scores. These ratings are used only to segment our analysis by acceptability. In Figure 7 we plot affinity scores for nouns in sentences with acceptability $\geq$ 4 (upon: 65, after: 73, by: 52, to: 54; total 244). We include the same plot but without acceptability filtering in the Appendix, Figure 18; lower acceptability sentences have lower affinity, reflecting that affinity is sensitive to linguistic acceptability.
In Figure 7, affinity scores for upon and after suggest that the distribution captures the form of the NPN: the model generally expects the two nouns to be the same. Moreover, the lower affinity for NPNs with by and to (and also the relative counts of acceptable generations) accords with prior characterization of the NPN: after and upon are more flexible in NPN use than other prepositions, and to is only semi-productive Jackendoff (2008).
As our objective is to test generalization to unseen NPNs, we generate a separate challenge dataset of $\sim 100$ nouns: Using the infinigram API Liu et al. (2024), we sample nouns that are not used in an NPN (with any of the four prepositions) in the Pile-train dataset (Gao et al. 2020; see E.2 for details). Though the affinity distribution for entirely unattested NPNs is more skewed toward lower affinities, we still see clear evidence of generalization (Figure 18).
@@ 6.2 Models generalize the comparative correlative’s category constraint
We test whether the distribution encodes the semantic category of the comparative adjective/adverb in the CC (see Weissweiler et al., 2022, for a recent study), which is of the form, e.g., The better your distribution, the more constructions it will encode. Whereas in §5.1, we showed that the in the CC has high global affinity, here we test whether the distribution also encodes an abstract slot constraint for the comparative adjective/adverb. Using the 54 CC examples from the CoGS dataset, we mask each comparative adjective/adverb, obtain the set of highest probability outputs at the masked position that sum to 98% probability mass, and calculate a comparative score: the percentage of this set that is a comparative adjective/adverb (see F).
Out of 99 comparative adjectives/adverbs, 95 score 100%; another 3 score $99\%$, and one (The higher up the nicer!) scores 86%, with non-comparatives ladder and mountain in the top 10 fills, perhaps because it is uncommon for a comparative to come at the end of the sentence. This result shows that the distribution captures the abstract syntactic (comparative adjective/adverb) constraint of the CC nearly perfectly, with 98/99 examples having scores $\geq 99\%$.
@@ 7 The Limits of Distributional Analysis
Figure 8: Green Day (a cxn) is present in top/bottom panels but without/with band, it has low/high global affinity (white = 0, black = 1).
Figure 9: The local affinity matrix reflects interactions (pink) between between band, Green, and Day, as expected.
@@
How far can we push these distributional approaches for identifying constructions in text? Could they allow us to identify a model’s complete constructicon bottom-up (see, e.g., Dunn, 2017, 2019, 2024; Feng et al., 2022; Lyngfelt et al., 2018; Xu et al., 2024)? We argue the answer is likely “no”: although affinity is a highly useful diagnostic of models’ knowledge of constructions, here we show that it cannot be directly equated with constructionhood.
First, consider the sentence Alice went to the hardware store and bought a hammer. The affinity matrix (Figure 7) reveals various interactions of interest. Some of these are likely constructionally mediated, including tense agreement in coordinated verb phrases (went, bought), as well as subject-verb (she, bought), head-modifier (hardware, store), and determiner-noun (a, hammer) dependencies. But other interactions are less obviously constructional, such as coreference (Alice, she) and semantic relatedness (hardware, hammer). So although the affinity matrix does reflect constructional relations, it also shows that other contextual interactions can produce high local affinities. Thus affinity is an insufficient criterion for constructionhood. [^5] Insufficiency holds for global affinity as well: Figure 19 shows that hardware has high global affinity (0.92).
Second, consider the sentence My favorite band is Green Day, which includes the non-compositional collocation Green Day (a well-known band name). Figure 10 shows that Green Day has low global affinity until an appropriate contextual trigger is given, i.e., band. This shows that even substantive constructions may exhibit low global affinities when the surrounding context is insufficient to trigger them. [^6] See also Figure 21 in the Appendix, for an additional example with the idiom kicked the bucket.
Additionally, Figure 10 shows the local affinity matrix for the Green Day example. As expected, it reflects affinity between band, Green, and Day. However, in a different context (I saw my favorite band, Green Day, in concert), the interactions between Green Day and band vanish (Figure 10). This appears to be due to the co-presence of an additional semantic cue (concert): since band and concert are both individually sufficient to cue the construction Green Day, neither individually exhibits strong pairwise affinity with Green Day (i.e., neither when masked substantially changes the model’s output distribution for Green or Day).
Taken together, the hardware store and Green Day examples show that affinity is neither sufficient nor necessary for a construction to be present. Insufficiency arises out of the tension between contextual and constructional interactions. Non-necessity arises from a methodological challenge: eliciting global affinity requires masking, but under mask, the context may not trigger a construction. These results thus warrant caution in mapping between observables (affinities) and hypothetical constructs (like constructions). They also suggest avenues for future work, which we explore below.
@@ 9 Conclusion
Our analysis is limited to a single (bidirectional) masked language model, RoBERTa, and could be rerun on other models of different sizes. In choosing masked language models, a straightforward analysis of the sort that we have performed is limited to words that are single tokens; multi-token generalization of these methods is left to future work. Our approach for §4.3 removed numerous examples, leaving relatively few for analysis; nonetheless the approach still recovered mislabeled examples. We studied only English constructions, and future work should look to apply these methods to other languages.
@@ Acknowledgements
Joshua Rozner was supported by the Institute for Computational and Mathematical Engineering at Stanford University. Leonie Weissweiler was supported by a postdoctoral fellowship from the German Research Foundation (DFG, WE 7627/1-1). Kyle Mahowald acknowledges funding from NSF CAREER grant 2339729. We thank Shijia Zhou for providing precise numeric results for prior work on the EAP, AAP, and CEC.
@@ Appendix A Supplement: Methods
Calculating a local affinity matrix is more expensive than calculating global affinities: $n+n^{2}$ forward passes versus $n$. Note that the affinity matrix is asymmetric—positions (i,j) and (j,i) involve three different distributions and are computed as JSD(A,B) vs. JSD(C,B)—and that the diagonal of the matrix, $a_{i,i}$ is 0.
We use RoBERTa large. Most experiments are run locally on an M3 Macbook Pro or on a single Nvidia RTX A6000 GPU on a cluster. Total time to run all experiment code is less than an hour.
@@ B.1 Data and preprocessing
B.2 Supplement for Section 4.1: Models distinguish the CEC from the EAP and AAP in their output distributions
Figure 11: This figure is the same as Figure 4, but without correction of mislabeled examples. As only 5 examples differ, the plots are very similar.
@@
B.3 Supplement for Section 4.2: Models capture causal relations in the CEC
@@ B.3.1 Dataset (multiple-that)
B.4 Supplement for Section 4.3: Affinity patterns distinguish the EAP and AAP
Figure 12: This plot shows which interactions, on average, most substantially differ between EAP and AAP examples. From highest to lowest in absolute value we have that with itself, so with adj, so with itself, verb with itself, and so with verb.
Figure 13: This figure is the same as Figure 4 but with labels for examples which would likely cluster with the wrong class or are on the boundary.
Figure 14: Multiplot version of Figure 4. UMAP projections for EAP (orange) and AAP (blue). Each row corresponds to the number of dimensions that are used for the projection (2-6). Columns correspond to different random seeds. Potentially “misclassified” sentences (those that were near the class boundary) are labeled with their original dataset IDs for discussion.
As described in §4.3, we identify potentially salient differences in the EAP and AAP patterns: we compute a single average (position-wise) affinity matrix for all EAP examples and another average matrix for all AAP examples. We fill the diagonal of the matrices with the average global affinity score (as probability). We then subtract these two average matrices and then take the position-wise absolute value. The largest values in the resulting matrix provide the interactions that seem to most distinguish the EAP and AAP. The heatmap in Figure 12 shows the most potentially informative interactions across EAP and AAP examples. Here we see that the most different interactions (between EAP and AAP examples) on average are that with itself, so with adj, and so with itself.
After running the pre-processing pipeline for tokenization we have 68 AAP and 24 EAP examples (see B.1). For our analysis, we exclude any that have multitokenized words in one of the seven positions we consider, or for which Spacy Honnibal et al. (2020) fails to label a POS. This leaves us with 14 EAP and 26 AAP examples for the analysis.
All UMAP projections use n_neighbors 10 and min_dist 0.1. Figure 14 shows UMAP plots using different numbers of affinities: each row corresponds to the number of informative dimensions that are chosen using the heatmap. The two columns show two random seeds for the UMAP projection. From the multiplot, we see that using five dimensions produces seems to mostly separate the two classes, and we use five dimensions in the main text (Figure 4).
Finally, we label five sentences in the plots (Figure 14) which either look to violate separability or which are on the boundary.
@@
    Label 10: It’s his lucky quarter and Pop feels so bad that Lucky lost it.
    After our CEC correction (relabeled to AAP from CEC), this would likely be misclassified using UMAP, since it clusters with EAP.
@@
    Label 99: An hour later, however, they’re still alive which confuses Elijah and Rebekah, as they were so positive that Klaus originated their bloodline and were sure it wasn’t Kol Mikaelson (Nathaniel Buzolic).
    This is epistemic, and could be misclassified using UMAP.
@@
    Label 113, In the police court, Mrs. Jones says she was so shocked that her husband had the box.
    This is affective and would likely be misclassified using UMAP.
@@
    Label 119: I am so sure that the lack of men on stage made some men feel excluded.
    This is epistemic, and could be misclassified using UMAP.
@@
    Label 301: I am so optimistic that I made the best choice.
    This is labeled as epistemic, though optimism conveys some degree of affect. This label is arguably ambiguous.
Appendix C Supplement for CoGS (Section 5.1)
CoGs has the following counts of examples for each construction type:
@@
    Causative-with: 50 (for with)
@@
    Comparative correlative: 54 (for the)
@@
    Much-less: 50 (for much and less)
@@
    Way-manner: 54 (for way)
We did not have any errors, so the number of examples reported in Figure 7 are exactly these, except for the comparative correlative “the”, for which there are $2\times 54=108$, since we treat the two the as a single class in the analysis.
Appendix D Supplement for MAGPIE (Section 5.2)
D.1 Data Sample